# Centralized Reward Agent for Knowledge Sharing and Transfer in Multi-Task Reinforcement Learning

### v2

@@ 6 Discussion and Conclusion
We propose a novel framework named CenRA that innovatively integrates reward shaping into MTRL. The framework not only shares domain knowledge across tasks to improve learning efficiency but also effectively addresses the sparse-reward challenge. Specifically, the centralized reward agent functions as a knowledge pool, responsible for knowledge distillation and distribution across multiple tasks. Furthermore, our information synchronization mechanism mitigates knowledge distribution disparities across tasks, ensuring optimal system-wide performance. Experiments demonstrate that the dense knowledge rewards generated by the CRA efficiently guide the policy agents’ learning, leading to faster convergence compared to baseline methods. The CenRA also performs superiorly and robustly in transferring acquired knowledge to new tasks.
CenRA’s current limitation is its requirement for uniform state and action shapes across tasks. Future improvements could explore preprocessing techniques to adapt the framework for varying state and action shapes to broaden CenRA’s applicability. Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal. A more flexible approach, potentially through adaptive regulation of these weights, could further enhance the framework. Besides, the performance weight might favor underperforming tasks, achieving overall balance, but potentially limiting the upper bound of top-performing tasks. This issue could necessitate a better mechanism for trade-offs.
@@ Appendix A Mutli-Task Experimental Configurations
We conduct experiments in four domains with multiple tasks: Meta-World, 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 8. The Meta-World tasks illustration is adapted from (Yu et al., 2020b, ).
@@
Appendix B Experimental Configurations
@@ B.1 Network Structures
B.2 Computing Resources
@@
B.3 Hyperparameters
Table 5: The hyperparameters of CenRA used in our experiments.
Module Hyperparameters Values
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $5000$
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 128
burn-in steps $10000$
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $10000$

### v3

@@ 6 Discussion and Conclusion
We propose a novel framework CenRA that integrates reward shaping into multi-task reinforcement learning. The framework shares domain knowledge across tasks to improve learning efficiency and effectively addresses the sparse-reward challenge. Specifically, the centralized reward agent (CRA) functions as a knowledge pool, responsible for distilling and distributing knowledge across tasks. Furthermore, the information synchronization mechanism mitigates imbalances in knowledge distribution, ensuring optimal system-wide performance. Experiments demonstrate that dense knowledge rewards generated by the CRA effectively guide policy learning, leading to faster convergence than baseline methods. CenRA also demonstrates superior and robust transferability to new tasks.
CenRA’s main limitation is its requirement for consistent state and action dimensions across tasks. Future work could explore preprocessing techniques to adapt the framework to varying task structures, broadening its applicability. Additionally, the fixed trade-off between similarity weight and performance weight may not be ideal. A more flexible approach, such as adaptive weight regulation, could further enhance the framework. Moreover, the performance weight might favor underperforming tasks to achieve overall balance, but could limit the performance ceiling of high-performing tasks, indicating the need for a more effective trade-off mechanism.
Acknowledgments and Disclosure of Funding
This work was supported by an Academic Research Grant MOE-T1 251RES2408 and a Research Scholarship from the Ministry of Education, Singapore.
@@ Appendix A Mutli-Task Experimental Configurations
We conduct experiments in four domains with multiple tasks: Meta-World, 2DMaze, 3DPickup, and MujocoCar. The detailed configurations of each task are illustrated in Figure 5. The Meta-World tasks illustration is adapted from [Yu et al., 2020b].
@@
Appendix B Network Structures and Hyperparameters
@@ B.1 Network Structures
B.2 Hyperparameters
Table 4: The hyperparameters of CenRA used in our experiments.
Module Hyperparameters Values
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
policy networks update frequency (steps) 22
target networks update frequency (steps) 11
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps 50005000
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 128
burn-in steps 1000010000
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}_{i}|$ $1\times 10^{6}$
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
policy networks update frequency (steps) 22
target networks update frequency (steps) 11
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps 1000010000
B.3 Computing Resources
@@
Appendix C What Has the Centralized Reward Agent Learned?
(a) Maze #1
(b) Maze #2
(c) Maze #3
(d) Maze #4
Figure 7: The actions yielding the maximum knowledge rewards in the four 2DMaze tasks.